# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ IV.1.2 Second-order interference of one superbunching light beam
Figure 14 shows the experimental scheme to observe two-photon superbunching with $2N$ different paths. Single-mode continuous-wave laser is employed as the light source. After passing through pinhole P1, the light is scattered by a rotating groundglass, RG1. Then the scattered light is filtered with another pinhole, P2. The size of P2 should be less than the coherence area of scattered light at P2 plane. Then the filtered light is scattered by RG2. Repeating the process $N$ times and superbunching pseudothermal light with $g^{(2)}(0)$ equaling $2^{N}$ can be generated.
@@
Figure 15: Normalized second-order coherence functions with one (a), two (b), and three (c) rotating groundglasses Bai et al. (2017). $t_{1}-t_{2}$ is the time difference between two photon detection events at D1 and D2. $g^{(2)}(t_{1}-t_{2})$ is the normalized second-order coherence function in Glauber’s quantum optical coherence theory, which is equivalent to the normalized two-photon probability function.
@@
Figure 15(a)-(c) shows the measured temporal second-order coherence function of light with one, two, and three rotating groundglasses, respectively Bai et al. (2017). $g^{(2)}(0)$ equals $1.96\pm 0.01$ in Fig. 15(a), in which there is one rotating groundglass and two-photon bunching is observed. The reason why it is less than 2 is due to part of the un-scattered laser light contribute to the two-photon coincidence count. The source with one rotating groundglass is pseudothermal light source invented by Martienssen and Spiller Martienssen and Spiller (1964). $g^{(2)}(0)$ equals $3.80\pm 0.04$ in Fig. 15(b) when there are two rotating groundglass. When there are three rotating groundglass, $g^{(2)}(0)$ equals $7.10\pm 0.07$ in Fig. 15(c). Two-photon superbunching is observed in Figs. 15(b) and 15(c).
@@
Another method to increase the degree of second-order coherence of light is the two-photon probability amplitudes changes with time. Figure 16 shows the scheme for observing two-photon superbunching by changing the two-photon probability amplitudes. The light source shown in Fig. 16 as the light source in a HBT interferometer shown in Fig. 11. There are two different paths for photons a and b to trigger a two-photon coincidence count in a HBT interferometer. The two-photon probability amplitudes are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$.
@@
Figure 16: Simplified superbunching pseudothermal light sourc Zhou et al. (2019b). IM is intensity modulator and RG is rotating groudglass.
@@
The difference between usual pseudothermal light source and the one shown in Fig. 16in the HBT interferometer is $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ will change with time in the latter due to that intensity modulator (IM) is employed to modulate the intensity of incidental laser light. When these two different paths to trigger a two-photon coincidence count are indistinguishable, the temporal two-photon probability function
@@
where $\gamma(t_{1}-t_{2})=\langle I(t_{1})I(t_{2})\rangle/(\langle I(t_{1})\rangle\langle I(t_{2})\rangle)$ is the correlation function of laser light introduced by IM, $I(t_{1})$ and $I(t_{2})$ are the intensities of light before RG, respectively. When $\gamma(t_{1}-t_{2})$ equals 1, Eq. (IV.1.2) describes normal two-photon bunching of thermal light. When $\gamma(t_{1}-t_{2})$ is larger than 1, two-photon suerpbunching can be observed.
Figure 17: Observed two-photon superbunching with changing probability amplitudes Zhou et al. (2019b).
Figure 17 shows the experimental measured temporal second-order coherence function of superbunching pseudotheral light with the scheme shown in Fig. 16. Two-photon superbuching with $g^{(2)}(0)$ equaling 20.45 is observed Zhou et al. (2019b).
@@ IV.1.3 Second-order interference of one laser light beam
Assuming S is a single-mode continuous-wave laser in Fig. 11. Same method can be employed to calculate two-photon probability function of laser light in a HBT interferometer. There are two different paths for photons a and b to trigger a two-photon coincidence count. The probability amplitudes for these two paths are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$. If these two different ways are indistinguishable, the two-photon probability is
@@
$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
It seems that two-photon bunching should also be observed for single-mode continuous-wave laser light in a HBT interferometer. However, it was proved that photon detection events are independent when a single-mode continuous-wave laser light beam in a HBT interferometer Arecchi (1965); Freed and Haus (1965). It is predicted in Glauber’s quantum optical coherence theory that the degree of second-order coherence of coherent state equals 1 Glauber (1963a), which indicates that photon detection events are independent in single-mode continuous-wave laser light.
The reason why there is no two-photon bunching of single-mode continuous-wave laser light lies at the difference between the superposition principles in classical and quantum theories. In classical theory, assuming there are two fields at space-time coordinate $(\vec{r},t)$, the final field is the superposition of these two fields,
@@
If $\vec{E}_{1}(\vec{r},t)$ is identical to $\vec{E}_{2}(\vec{r},t)$, the superposition will give $2\vec{E}_{1}(\vec{r},t)$. The intensity at that point will be four times of the one of $\vec{E}_{1}(\vec{r},t)$.
This is not true for the superposition in quantum theory. Assuming there are two wave functions at space-time coordinate $(\vec{r},t)$, the final wave function is the superposition of these two,
@@
If $\Psi_{1}(\vec{r},t)$ is identical to $\Psi_{2}(\vec{r},t)$, the superposition will not give $2\Psi_{1}(\vec{r},t)$. The results of superposition of two identical states equals one of the state, not two times of the individual state Bohm (2012). The wave function can be related to the probability amplitude Feynman et al. (2010),
@@
The superposition of two probability amplitudes for two identical paths will not give two times of each probability amplitude, but only one probability amplitude. The two different paths in Eq. (65) are identical for single-mode continuous-wave laser light. The two-photon probability function is
@@
no second-order interference exists and $A(S,D_{1};S,D_{2})$ is the probability amplitude for two photons emitted by S going to D1 and D2, respectively.
Comparing the two-photon probability functions of thermal and laser light in a HBT interferometer, the reason why these two cases are different is due to the initial phases of photons in thermal light are random and the initial phases of photons in laser light are identical within one coherence volume. This difference makes two different paths, $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ different for thermal light and identical for laser light. Hence two-photon bunching can be observed with thermal light and can not be observed for single-mode continuous-wave laser light.
@@ IV.2 Second-order interference of two light beams
The first second-order interference of two light beams experiment are correlation measurement of photon pairs generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). Pfleegor and Mandel employed the second-order correlation of light to prove there exist transient first-order interference pattern for two independent lasers at low light level Pfleegor and Mandel (1967). The most famous second-order interference of two light beams is the second-order interference of entangled photon pairs in a HOM interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). Many other second-order interference of light experiments were performed, please see Refs. Belinski and Klyshko (1993); Paul (1986); Ou (2007); Shih (2020) for details. This section will concentrate on applying Feynman’s path integral to discuss the second-order interference of two light beams in a HOM interferometer.
Figure 18: Second-order interference of two light beams in a HOM interferometer. S1 and S2 are two light sources. BS is a 1:1 nonpolarizing beam splitter. D1 and D2 are two photo detectors. CC is coincidence count detection system.
The second-order interference of two light beams in a HOM interferometer is shown in Fig. 18. Two light beams emitted by source S1 and S2 are incident to two input ports of a 1:1 beam splitter (BS). The output light beams from BS are detected by two single-photon detectors, D1 and D2. The output of these two detectors are sent into a two-photon coincidence count detection system to measure the second-order correlation.
@@ IV.2.1 Second-order interference of entangled photon pairs
Assume two photons in the same entangled photon pair are in the scheme shown in Fig. 18, which is similar as the original second-order interference with entangled photon pairs Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). There are two different paths for photons a and b to trigger a two-photon coincidence count in D1 and D2. For simplicity, assume photon a is emitted by S1 and photon b is emitted by S2. The first path is photon a is detected by D1 and photon b is detected by D2. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{1};b,D_{2}}$. The second path is photon a is detected by D2 and photon b is detected by D1. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{2};b,D_{1}}$. If these two different paths are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ $=$ $\langle|A_{a,D_{1};b,D_{2}}+A_{a,D_{2};b,D_{1}}|^{2}\rangle.$ (70)
There is an extra $\pi/2$ phase of the reflected photon compared to the transmitted photon introduced by BS Loudon (2000), which can not be ignored in HOM interferometer as the one in HBT interferometer. Equation (70) can be simplified as
@@
$=$ $\langle|e^{i\varphi_{a}}K_{aD_{1}}e^{i\varphi_{b}}K_{bD_{2}}+e^{i\varphi_{a}}e^{i\pi/2}K_{aD_{2}}e^{i\varphi_{b}}e^{i\pi/2}K_{bD_{1}}|^{2}\rangle$
@@
where $\varphi_{a}$ and $\varphi_{b}$ are the initial phases of photon a and b, respectively, $K_{aD_{1}}$ is the photon’s Feynman propagator for photon a emitted by S1 goes to D1 and the meanings of other symbols are defined similarly.
Since $\varphi_{a}$ and $\varphi_{b}$ are random for entangled photon pairs, the probability to detect one photon at D1 or D2 is a constant. The normalized two-photon probability function is proportional to the two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$.
Both temporal and spatial second-order interference can be calculated. First calculate the temporal case. Assume D1 and D2 are in the symmetrical positions about BS, all the properties such as polarization, frequency are the same for photons a and b. With the same method as the one in Sect. IV, Eq. (71) can be simplified as
@@
Figure 19: HOM dip for entangled photon pairs with different frequency bandwidths. 1, 2, and 3 correspond to $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
When the time difference between these two photon detection events is zero, $P^{(2)}(t_{1}-t_{2})$ equals zero, which means it is impossible to detect two photons at D1 and D2 simultaneously. These two photons are in the same output port of BS. When $|t_{1}-t_{2}|$ increases, $P^{(2)}(t_{1}-t_{2})$ increases from zero to one. $P^{(2)}(t_{1}-t_{2})$ equals 1 when $|t_{1}-t_{2}|$ is much longer than the coherence time, which indicates the detections of these two photons are independent.
Similar method can be employed to calculate spatial two-photon probability function
@@
where $x_{1}$ and $x_{2}$ are the transverse coordinates of D1 and D2, respectively, $d$ is the size of one dimension source. Paraxial approximation and one-dimension are assumed to simplify the results. Experiments with entangled photon pairs confirmed the result in Eq. (73) Kim et al. (2006); Lee and van Exter (2006); Devaux et al. (2019).
It is interesting to comparing two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pair in a HOM interferometer. Two-photon bunching of thermal light shown in Fig. 12 looks like the contrary effect of HOM dip shown in Fig. 19. Photons have the tendency to bunch together when these two detectors are at symmetrical positions in Fig. 12 while photons have the tendency to avoid each other when two detectors are at symmetrical positions in Fig. 19. Even through different phenomena are observed with types of light in different interferometers, the physics of these two phenomena are the same. As shown in Eqs. (IV.1.1) and (71), they are results of two-photon interference. The only difference is that two-photon bunching of thermal light is caused by constructive two-photon interference and HOM dip with entangled photon pairs is caused by destructive two-photon interference.
Entangled photon pairs are not necessary to observe HOM dip shown in Fig. 19. The same HOM dip can be observed with two independent and identical single-photon states as the input of HOM interferometer. The theoretical calculations are exactly the same as the one for entangled photon pairs and the results are experimentally verified by several groups Legero et al. (2004); Beugnon et al. (2006); Flagg et al. (2010); Qian et al. (2016).

### v2

@@ IV.1.2 Second-order interference of one superbunching light beam
Figure 14 shows the experimental scheme to observe two-photon superbunching with $2N$ different paths. A single-mode continuous-wave laser is employed as the light source. After passing through pinhole P1, the light is scattered by a rotating groundglass, RG1. Then the scattered light is filtered by another pinhole, P2. The size of P2 should be less than the coherence area of the scattered light at P2 plane. Then the filtered light is scattered by RG2. Repeating the process $N$ times and superbunching pseudothermal light with $g^{(2)}(0)$ equaling $2^{N}$ can be generated.
@@
Figure 15: Normalized second-order coherence functions with one (a), two (b), and three (c) rotating groundglasses Bai et al. (2017). $t_{1}-t_{2}$ is the time difference between two photon detection events at D1 and D2. $g^{(2)}(t_{1}-t_{2})$ is the normalized second-order coherence function in Glauber’s quantum optical coherence theory, which is equivalent to the normalized two-photon probability function in quantum optical coherence theory based on Feynman’s path integral.
@@
Figure 15(a)-(c) shows the measured temporal second-order coherence functions of light with one, two, and three rotating groundglasses, respectively Bai et al. (2017). $g^{(2)}(0)$ equals $1.96\pm 0.01$ in Fig. 15(a), in which there is one rotating groundglass and two-photon bunching is observed. The reason why it is less than 2 is due to that part of the un-scattered laser light contributes to the two-photon coincidence count. The source with one rotating groundglass is pseudothermal light source invented by Martienssen and Spiller Martienssen and Spiller (1964). $g^{(2)}(0)$ equals $3.80\pm 0.04$ in Fig. 15(b) when there are two rotating groundglasses. When there are three rotating groundglasses, $g^{(2)}(0)$ equals $7.10\pm 0.07$ in Fig. 15(c). Two-photon superbunching is observed in Figs. 15(b) and 15(c).
@@
Another method to increase the degree of second-order coherence of light is that the two-photon probability amplitudes changes with time. Figure 16 shows the scheme for observing two-photon superbunching by changing the two-photon probability amplitudes. The light source shown in Fig. 16 is the light source in a HBT interferometer shown in Fig. 11. There are two different paths for photons a and b to trigger a two-photon coincidence count in a HBT interferometer. The two-photon probability amplitudes are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$.
@@
Figure 16: Simplified superbunching pseudothermal light source Zhou et al. (2019b). IM is intensity modulator and RG is rotating groundglass.
@@
The difference between usual pseudothermal light source and the light source shown in Fig. 16 in a HBT interferometer is that $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ will change with time in the latter case. An intensity modulator (IM) is employed to modulate the intensity of incidental laser light. When these two different paths to trigger a two-photon coincidence count are indistinguishable, the temporal two-photon probability function is
@@
where $\gamma(t_{1}-t_{2})=\langle I(t_{1})I(t_{2})\rangle/(\langle I(t_{1})\rangle\langle I(t_{2})\rangle)$ is the correlation function of laser light introduced by IM, $I(t_{1})$ and $I(t_{2})$ are the intensities of light at time $t_{1}$ and $t_{2}$ before RG, respectively. When $\gamma(t_{1}-t_{2})$ equals 1, Eq. (IV.1.2) describes normal two-photon bunching of thermal light. When $\gamma(t_{1}-t_{2})$ is larger than 1, two-photon superbunching can be observed.
Figure 17: Two-photon superbunching with changing probability amplitudes Zhou et al. (2019b).
Figure 17 shows the experimental measured temporal second-order coherence function of superbunching pseudothermal light with the scheme shown in Fig. 16. Two-photon superbunching with $g^{(2)}(0)$ equaling 20.45 is observed Zhou et al. (2019b).
@@ IV.1.3 Second-order interference of one laser light beam
Assume S is a single-mode continuous-wave laser in Fig. 11. Same method can be employed to calculate two-photon probability function of laser light in a HBT interferometer. There are two different paths for photons a and b to trigger a two-photon coincidence count. The probability amplitudes for these two paths are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$. If these two different paths are indistinguishable, the two-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{2}}[A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})]|^{2}\rangle.$
It seems that two-photon bunching should be observed for single-mode continuous-wave laser light in a HBT interferometer. However, it was proved that photon detection events are independent for a single-mode continuous-wave laser light beam in a HBT interferometer Arecchi (1965); Freed and Haus (1965). It is predicted that the degree of second-order coherence of coherent state equals 1 in Glauber’s quantum optical coherence theory Glauber (1963a), which indicates that the photon detection events are independent in single-mode continuous-wave laser light.
The reason why there is no two-photon bunching of single-mode continuous-wave laser light lies at the difference between the superposition principles in classical and quantum theories. In classical theory, if there are two fields at space-time coordinate $(\vec{r},t)$, the final field is the superposition of these two fields,
@@
If $\vec{E}_{1}(\vec{r},t)$ is identical to $\vec{E}_{2}(\vec{r},t)$, the superposition will result in $2\vec{E}_{1}(\vec{r},t)$. The intensity at the point will be four times of the one of $\vec{E}_{1}(\vec{r},t)$.
This is not the case for the superposition principle in quantum theory. Assuming there are two wave functions at space-time coordinate $(\vec{r},t)$, the final wave function is the superposition of these two wave functions,
@@
If $\Psi_{1}(\vec{r},t)$ is identical to $\Psi_{2}(\vec{r},t)$, the superposition will not give $2\Psi_{1}(\vec{r},t)$. The result of superposition of two identical states equals the individual state, not two times of the individual state Bohm (2012). The wave function can be related to the probability amplitude Feynman et al. (2010),
@@
The superposition of two probability amplitudes for two identical paths will not get two times of each probability amplitude, but only one probability amplitude. The two different paths in Eq. (65) are identical for single-mode continuous-wave laser light. The two-photon probability function is
@@
no second-order interference pattern exists and $A(S,D_{1};S,D_{2})$ is the probability amplitude for two photons emitted by S going to D1 and D2, respectively.
Comparing the two-photon probability functions of thermal and laser light in a HBT interferometer, the reason why these two cases are different is due to the initial phases of photons are random in thermal light and the initial phases of photons are identical in laser light within one coherence volume. This difference makes two different paths, $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$, different for thermal light and identical for laser light. There is two-photon bunching for thermal light and no two-photon bunching for single-mode continuous-wave laser light.
@@ IV.2 Second-order interference of two light beams
The first second-order interference experiment with two light beams is correlation measurement of photon pairs generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). Pfleegor and Mandel employed the second-order correlation of light to prove that there exists transient first-order interference pattern for two independent lasers at low light level Pfleegor and Mandel (1967). The most famous second-order interference of two light beams is the second-order interference of entangled photon pairs in a HOM interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). Many other second-order interference experiments with light were performed, please see Refs. Belinski and Klyshko (1993); Paul (1986); Ou (2007); Shih (2020) for details. This section will concentrate on applying Feynman’s path integral to discuss the second-order interference of two light beams in a HOM interferometer.
Figure 18: Second-order interference of two light beams in a HOM interferometer. S1 and S2 are two light sources. BS is a 1:1 non-polarizing beam splitter. D1 and D2 are two single-photon detectors. CC is a two-photon coincidence count detection system.
The scheme for the second-order interference of two light beams in a HOM interferometer is shown in Fig. 18. Two light beams emitted by two sources, S1 and S2, are incident into two input ports of a 1:1 non-polarizing beam splitter, BS. The output light beams from BS are detected by two single-photon detectors, D1 and D2. The output of these two detectors are sent into a two-photon coincidence count detection system to measure the second-order correlation.
@@ IV.2.1 Second-order interference of entangled photon pairs
Assume two photons in one entangled photon pair are in the scheme shown in Fig. 18, which is similar as the original second-order interference of entangled photon pairs Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). There are two different paths for photons a and b to trigger a two-photon coincidence count in D1 and D2. For simplicity, assume photon a is emitted by S1 and photon b is emitted by S2. The first path is that photon a is detected by D1 and photon b is detected by D2. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{1};b,D_{2}}$. The second path is that photon a is detected by D2 and photon b is detected by D1. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{2};b,D_{1}}$. If these two different paths are indistinguishable, the two-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{2}}[A_{a,D_{1};b,D_{2}}+A_{a,D_{2};b,D_{1}}]|^{2}\rangle.$
There is an extra $\pi/2$ phase for the reflected photon compared to the transmitted photon introduced by BS Loudon (2000), which can not be ignored in HOM interferometer. Equation (70) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (71)
$\propto$ $\langle|e^{i\varphi_{a}}K_{aD_{1}}e^{i\varphi_{b}}K_{bD_{2}}+e^{i\varphi_{a}}e^{i\pi/2}K_{aD_{2}}e^{i\varphi_{b}}e^{i\pi/2}K_{bD_{1}}|^{2}\rangle$
@@
where $\varphi_{a}$ and $\varphi_{b}$ are the initial phases of photons a and b, respectively, $K_{aD_{1}}$ is the photon’s Feynman propagator for photon a emitted by S1 goes to D1 and the meanings of other symbols are defined similarly. $\varphi_{a}$ and $\varphi_{b}$ are random for entangled photon pairs, the probability to detect one photon at D1 or D2 is a constant. The normalized two-photon probability function is proportional to the two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$.
Both temporal and spatial second-order interference patterns can be calculated. Let us first calculate the temporal case. Assume D1 and D2 are in the symmetrical positions of the BS. All the properties such as polarization, frequency, are the same for photons a and b. With the same method as the one in Sect. IV, Eq. (71) can be simplified as
@@
Figure 19: Simulated HOM dip for entangled photon pairs with different frequency bandwidths. 1, 2, and 3 correspond to $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
When the time difference between these two photon detection events is zero, $P^{(2)}(t_{1}-t_{2})$ equals zero, which means that it is impossible to detect two photons at D1 and D2 simultaneously. These two photons are in the same output port of BS. When $|t_{1}-t_{2}|$ increases, $P^{(2)}(t_{1}-t_{2})$ increases from zero to one. $P^{(2)}(t_{1}-t_{2})$ equals 1 when $|t_{1}-t_{2}|$ is much longer than the coherence time, which indicates that the detections of these two photons are independent.
Similar method can be employed to calculate the spatial two-photon probability function,
@@
where $x_{1}$ and $x_{2}$ are the transverse coordinates of D1 and D2, respectively, $d$ is the size of one dimension source. Paraxial approximation and one-dimension case are assumed to simplify the results. Experiments with entangled photon pairs confirmed the result in Eq. (73) Kim et al. (2006); Lee and van Exter (2006); Devaux et al. (2019).
It is interesting to compare two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pair in a HOM interferometer. Two-photon bunching of thermal light shown in Fig. 12 looks like a contrary effect of HOM dip shown in Fig. 19. Photons have the tendency to bunch together when these two detectors are at symmetrical positions in Fig. 12. Photons have the tendency to leave the BS from the same output port when these two detectors are at symmetrical positions in Fig. 19. Even through different phenomena are observed with different types of light in different interferometers, the physics of these two phenomena are the same. As shown in Eqs. (IV.1.1) and (71), they are results of two-photon interference. The only difference is that two-photon bunching of thermal light is caused by constructive two-photon interference and HOM dip with entangled photon pairs is caused by destructive two-photon interference.
Entangled photon pairs are not necessary to observe HOM dip shown in Fig. 19. The same HOM dip with 100% visibility can be observed with two independent and identical single-photon states. Theoretical calculations are exactly the same as the one for entangled photon pairs and the results are experimentally verified by several groups Legero et al. (2004); Beugnon et al. (2006); Flagg et al. (2010); Qian et al. (2016).